OpenStudy (anonymous):

what is the value of (i)to the power of 4n+1

jhonyy9 (jhonyy9):

do you know the power of i ? so like i^2 = -1 and i^3 = (-1)*i = -i and hope you can continue it so you need to see that when i has a power even what is a power of 2 so than this will result what in second step ? so because i^2 = -1 so this mean that than you continue this line with exponent what is a power of 2 so than for the power of i will get i^4 = (-1)^2 = 1 so after this will get 1 on power of power of 2 what will result in continue 1 so in case of this exponent from this presently your exercise how you think what will result ?